# Dynamic Approximate Maximum Matching in the Distributed Vertex Partition Model

### v2

@@ 2. Preliminaries
Lemma 1 (Algorithm Spreading (Li et al., 2024)).
@@
Lemma 4 (Data Processing Inequality, see Theorem 2.8.1 in (Cover, 1999)).
@@ 3. A Fully Dynamic Approximate Maximum Matching Algorithm
In this section, we give an algorithm for handling single edge updates.
@@
We first describe how to handle the insertion of a single edge $\{u,v\}$. Players $P_{u}$ and $P_{v}$ check whether $u$ and $v$ were both matched, in which case no further action is required. Similarly, if neither of them was matched, players $P_{u}$ and $P_{v}$ simply add the edge $\{u,v\}$ to the matching. Note that this cannot cause any 3-augmenting paths, because this would require both $u$ and $v$ to have a free neighbor, contradicting the assumption that the matching prior to the update was maximal. Finally, consider the case that only one of the two nodes, say $v$, was matched to some node $w$ whereas $u$ is free. In that case, $P_{v}$ instructs $P_{w}$ to check with all neighbors of $w$ whether $w$ has any free neighbor. This requires $P_{w}$ to send a single message to all players, and each player who has at least one free neighbor $x$ of $w$ (note that this is part of the players’ initial knowledge) responds affirmatively by sending $x$’s ID. If no such $x$ is found, we are again done, as it is straightforward to see that there cannot exist a 3-augmenting path. Otherwise, the four players $P_{u}$, $P_{v}$, $P_{w}$, and $P_{x}$ then coordinate to update the matching by replacing the edge $\{v,w\}$ with the two edges $\{u,v\}$ and $\{w,x\}$, which can be done in $O(1)$ additional rounds. Note that this results in all four involved vertices to become matched, and hence this cannot create any new 3-augmenting path in the resulting matching.
@@
      Suppose that $u$’s degree $d(u)$ is such that $d(u)\leqslant 2\sqrt{m}$. Then, $P_{u}$ uses Spreading (see Lem. 1) to broadcast the IDs of $u$’s neighbors, denoted by $N(u)$, to all players, which takes at most $O(\lceil\frac{\sqrt{m}}{k\beta}\rceil)$ rounds. As a result, every player $P_{i}$, who hosts a vertex $w_{i}$ that is matched to some $v_{i}\in N(u)$, knows that $v_{i}$ is matched to a neighbor of $u$. Next, $P_{i}$ checks with the other players whether $w_{i}$ has a free neighbor. We point out that $P_{i}$ may be hosting multiple vertices that are matched to (distinct) neighbors of $u$’s. In that case, $P_{i}$ performs this check via Spreading. By assumption, there are at most $2\sqrt{m}$ such matched vertices for which the players need to perform this check (in parallel), and hence Spreading takes $O(\lceil\frac{\sqrt{m}}{k\beta}\rceil)$ rounds to complete according to Lemma 1. Any player who hosts at least one free neighbor $x_{i}$ of $w_{i}$, picks one of them arbitrarily, and then broadcasts the ID of $x_{i}$; again, all of these (at most $2\sqrt{m}$) IDs are sent by the players via Spreading. Consequently, player $P_{u}$ learns at least one 3-augmenting path that $u$ is part of (if one exists). $P_{u}$ chooses one of these paths, say $(u,v_{i},w_{i},x_{i})$, and then directly coordinates with players $P_{v}$, $P_{w_{i}}$, and $P_{x_{i}}$ to replace the edge $\{v_{i},w_{i}\}$ with the edges $\{u,v_{i}\}$ and $\{w_{i},x_{i}\}$ in the matching. As this does not cause any vertices to become free that were matched before the edge deletion, this operation cannot create new 3-augmenting paths.
@@
      Now, suppose that $d(u)>2\sqrt{m}$: Let $Z$ denote the set of nodes that are matched to $u$’s neighbors. By a simple counting argument, it follows that there is a subset $Z^{\prime}\subseteq Z$ of size at least $\frac{|Z|}{2}$, such that every $x\in Z^{\prime}$ has $d(x)\leqslant 2\sqrt{m}$. We describe how $P_{u}$ can contact the players of $\Gamma=\Theta\left(\log n\right)$ randomly sampled vertices in $Z$: Note that $P_{u}$ may not know $Z$ (and hence cannot determine the players hosting the vertices in $Z$), but only knows the set of neighbors $N(u)$. To implement this step, $P_{u}$ randomly samples $\Gamma$ nodes in $N(u)$ and broadcasts the set of sampled node IDs, denoted by $S$, via Spreading to all players, which takes $O\left(\lceil\Gamma/\beta k\rceil\right)=O\left(\lceil\frac{\log n}{\beta k}\rceil\right)$ rounds. For each sampled $v\in S$ that is matched to some $w\in Z$, player $P_{v}$ asks $P_{w}$ to broadcast the ID of $w$ via Spreading. It follows that, with high probability, $P_{u}$ obtains the node ID of some low-degree vertex $w^{\prime}\in Z^{\prime}$ that is matched to some $v^{\prime}\in N(u)$. (To obtain a Las Vegas algorithm, we can simply repeat this step, if necessary.) Then, the players $P_{u}$, $P_{v^{\prime}}$, and $P_{w^{\prime}}$ add the edge $\{u,v^{\prime}\}$ to the matching and unmatch the edge $\{v^{\prime},w^{\prime}\}$ instead. As a result, there cannot exist any 3-augmenting path traversing $\{u,w\}$, as we assumed that $u$ does not have any free neighbors. However, this operation may have introduced 3-augmenting paths starting at the now-unmatched node $w^{\prime}$.
@@
Finally, we observe the players do not need to keep track of any state across updates apart from the matched edges, which ensures that algorithm is also memoryless. ∎
@@ 4. Incremental $\frac{2}{3}$-approximate Maximum Matching under Batch-Dynamic Updates
In this section, we consider an adaptive adversary who can insert a batch of edges at once. The correctness of our algorithm rests on the following classic result:
Lemma 1 ((Hopcroft and Karp, 1973)).
From Lemma 1, we can see that, for any $\frac{2}{3}$-approximate maximum matching, there is no augmenting path with length $3$, which exists if and only if there is a matched edge whose both endpoints have free neighbors. Throughout this section, we restrict ourselves to incremental updates, i.e., batches of edge insertions. The following theorem shows our main result for batch dynamic algorithms.
@@ 4.1. Description and Analysis of the Algorithm
Our algorithm splits the batch of $\ell$ new edges into mini-batches of $\mathbf{b}=\lfloor\sqrt{k\beta}\rfloor$ edges each, and then it processes these mini-batches sequentially, such that the resulting matching after each mini-batch is maximal and does not contain 3-augmenting paths. Thus, we focus on describing the algorithm for just a single mini-batch of edges $B$. Without loss of generality, we assume that no edges in $B$ connect two matched nodes, as the algorithm can simply ignore these edges.
We first introduce some notation. We use $M_{0}$ to denote matching before the algorithm performs any updates for mini-batch $B$. We say that an edge is new if it is part of $B$, and old otherwise. Similarly, if an edge was inserted between $u$ and a free node $v$, we say that $v$ is a new free neighbor of $u$. Note that $v$ may be a new free neighbor for one node and simultaneously an old free neighbor for another.
The main algorithm consists of three conceptual parts that we call phases.
@@
Initially, we want to ensure that all players know $B$, and thus we broadcast the entire set to all players using procedure Spreading (see Lem. 1). In Phase 1, we focus on a subgraph $G_{1}$ that is small enough (i.e., consists of $O(\sqrt{k\beta})$ edges) to be collected and processed efficiently by just a single player. The edges forming $G_{1}$ consist of several subsets that we describe next:
@@
  • We will also process certain triangles in Phase 1 that may be part of 3-augmenting paths: To this end, consider any triangle $t=\{w,u,c\}$, for some vertices $u$, $w$, and $c$. We say that $t$ is a harmful triangle, if the edge $\{w,u\}$ is matched, $c$ is free, and at least one of $w$ or $u$ has an additional incident new edge to a free neighbor. Note that this includes the case that at least one of the other two unmatched edges of the triangle (i.e., $\{w,c\}$ or $\{u,c\}$) are new. Also note that there exists at most one triangle $\{c,w,u\}$ for a matched edge $\{w,u\}\in I_{1}$ before edge insertions (otherwise, there would have been a 3-augmenting path prior to processing $B$). The harmful triangles also contain such triangles. On the other hand, we say that $t$ is a harmless triangle if $\{w,u\}$ is matched, $c$ is either matched or free, and there does not exist a 3-augmenting path traversing $\{w,u\}$. Let $T$ be the set of edges forming harmful triangles, whereby $T$ also includes all incident free edges.
@@
Our analysis crucially relies on the invariant that there are no $3$-augmenting paths in subgraph $G_{1}$, which we guarantee by sending all of these edges to player $P_{1}$. To this end, $P_{1}$ first computes a maximal matching $M^{\prime}$ and then iteratively removes any 3-augmenting path $p=(a,b,c,d)$, where $\{b,c\}$ is currently matched, by unmatching $\{b,c\}$ and instead matching $\{a,b\}$ and $\{c,d\}$. Subsequently, all other players update the state of their vertices and edges by learning the computed matching. Both of these communication steps are done by invoking Algorithm Spreading. We use $M_{1}$ to denote the updated matching on the whole graph.
@@
Assume towards a contradiction that there is an edge $e$ between two vertices in $V\cup X$. Since both endpoints of $e$ are free vertices at the start of Phase 2, and we do not unmatch any previously matched vertex in Phase 1 (see Lem. 4(a)), it follows that that the endpoints of $e$ were free also at the start of Phase 1. Thus, $e$ must be a new edge, as otherwise the matching that we started with was not maximal. Since $e\in F$, it must have been included in graph $G_{1}$, on which we have computed a maximal matching in Phase 1, which is a contradiction. ∎
@@
We first show the result for nodes in $W$. Assume towards a contradiction that $w_{i}\in W$ has a new free neighbor $v$. Recall that $w_{i}$ is matched to $u_{i}$ who also has a new free neighbor $v^{\prime}$. Moreover, by the definition of $I$, node $w_{i}$ also has an old free neighbor $x\in X$ and it holds that $x\neq v^{\prime}$. If $v=v^{\prime}$, then $(w_{i},u_{i},v^{\prime})$ forms a harmful triangle $t$, as there is a 3-augmenting path $p=(x,w_{i},u_{i},v^{\prime})$. By the definition of graph $G_{1}$ (see Lem. 3), we know that $p$ is part of $G_{1}$, contradicting Lemma 4(c). Now, suppose that $v\neq v^{\prime}$. It follows that there exists a 3-augmenting path $(v,w_{i},u_{i},v^{\prime})$ that is part of $\hat{I}_{1}\subseteq G_{1}$, because both $\{v,w_{i}\}$ as well as $\{u_{i},v\}$ are new free edges. Again, this contradicts Lemma 4(c).
Now, assume towards a contradiction that some node $u_{i}\in U$ has a free old neighbor $x$. Let $x^{\prime}$ be the free old neighbor of $w_{i}$. In the case that $x\neq x^{\prime}$, there exists a 3-augmenting path $p=(x^{\prime},w_{i},u_{i},x)$. Since we did not unmatch any vertices in Phase 1, it follows that $p$ also existed in the matching at the start of the update, which contradicts that the previous matching did not contain 3-augmenting paths. On the other hand, if $x=x^{\prime}$, we know that $u_{i}$ must also have a free new neighbor $v$. It follows that $(x,w_{i},u_{i})$ forms a harmful triangle, because there is a 3-augmenting path $(x,w_{i},u_{i},v)$, thus contradicting Lemma 4(c).
@@
Consider the matching $M_{1}$ resulting from Phase 1. Every 3-augmenting path in $G$ traverses a matched edge in $I$.
@@
Suppose that there is a 3-augmenting path $(a,b,c,d)$, where the matched edge $\{b,c\}\notin I$. It is not possible that both $a$ and $b$ have a new incident edge, because then $(a,b,c,d)$ would have been part of $\hat{I}_{1}\subseteq G_{1}$ and processed in Phase 1. Furthermore, it is also not possible that both $a$ and $b$ have an incident old free edge, as in that case, the matching that we started with prior to the update had a 3-augmenting path, contradicting our assumption. Thus, it follows that either $\{a,b\}$ or $\{c,d\}$ must be a new edge, which allows us to conclude that $\{b,c\}\in I$. ∎
@@
We use the edge set $I$ (see Def. 5) to define a new bipartite graph $G_{X,W}$, on which we will operate in Phase 2: The vertices of $G_{X,W}$ consist of the sets $X$ and $W$. We include all edges from $G$ in $G_{X,W}$ that are of the form $\{x,w\}$, where $x\in X$ and $w\in W$, which ensures that $G_{X,W}$ is bipartite. In particular, we omit all edges between nodes in $W$ (recall that there cannot be any edges at this point between nodes in $X$ as all of them are free, see Lemma 6.) Next, we make use of the result of (Li et al., 2024), namely that there exists an algorithm that maintains a maximal matching on a graph with a small update time, when the adaptive adversary deletes a batch edges. A useful consequence of their result is that it yields a fast algorithm for computing a maximal matching on bipartite graphs, as we show next:
@@
Without loss of generality, assume that $|A|\leqslant|B|$. Since each player knows which of its vertices are in $A$ and which ones are in $B$, we can simply aggregate these counts at player $P_{1}$ and ensure that every player learns that $s=|A|$. Then, for each node $a\in A$, the player $P_{a}$ locally adds a neighbor $a^{\prime}$ to $H$ and pretends that the edge $\{a,a^{\prime}\}$ is matched, which ensures that there is a maximal matching on this extended graph. Note that the resulting graph partition is still balanced. Then, we remove all $\{a,a^{\prime}\}$ edges, to which the players respond by executing the update algorithm of (Li et al., 2024). Since the number of deleted edges is $s$, Theorem 1.2 in (Li et al., 2024) tells us that this algorithm completes in $O\left(\frac{s\log(n)}{\sqrt{\beta k}}\right)$ rounds, as required. ∎
@@
In the final phase, we operate on the subgraph $G_{*}^{\prime}=G[X^{\prime}\cup W\cup U\cup V]$. By definition, every edge in $I$ has at least one incident new edge (at the start of Phase 2), and the same is true for the vertices in $V$. This means that the number of vertices in $G_{*}^{\prime}$ is $O(\sqrt{k\beta})$, and thus the number of edges (new or old) in $G_{*}^{\prime}$ is $O({k\beta})$. We send all edges of $G_{*}^{\prime}$ to player $P_{1}$ in $O(1)$ rounds for local processing via Spreading (see Lem. 1). Starting from matching $M_{1}$, the local computation at $P_{1}$ produces the final matching $M$, by processing each $\{w_{i},u_{i}\}\in I$ in sequence, thereby iteratively augmenting the matching obtained in Phase 1. Note that we cannot simply instruct $P_{1}$ to locally compute a maximum matching on $G_{*}$, as we require the resulting matching to have certain properties, which we formally state in Lemma 11 below.
@@
For each edge $\{w_{i},u_{i}\}\in I$ ($i\geqslant 1$), player $P_{1}$ checks whether there is a 3-augmenting path $p=(x_{i},w_{i},u_{i},v)$ in $G_{*}^{\prime}$, for any free $v\in V$, where $x_{i}\in X^{\prime}$ denotes the vertex to which $w_{i}$ was matched in $M_{X,W}$, if it exists. If there is such a path, then $P_{1}$ updates $M$ by replacing $\{w_{i},u_{i}\}$ with the edges $\{x_{i},w_{i}\}$ and $\{u_{i},v\}$, before moving on to edge $\{w_{i+1},u_{i+1}\}$.
@@
Next, player $P_{1}$ inspects every (currently) matched edge $\{w_{i},u_{i}\}\in I$ and checks whether there exists any 3-augmenting path for any free $x\in X^{\prime}$ and $v\in V$; note that it is not necessarily the case that $x=x_{i}$. After all edges in $I$ have been processed in the above manner, $P_{1}$ sends the final matching $M$ to every player via Spreading.

### v3

@@ 2. Preliminaries
Lemma 1 (Algorithm Spreading (li2024dynamic)).
@@
Lemma 4 (Data Processing Inequality, see Theorem 2.8.1 in (cover1999elements)).
@@ 3. A Fully Dynamic Approximate Maximum Matching Algorithm
In this section, we give an algorithm for handling single edge updates (Result 1).
@@
We first describe how to handle the insertion of a single edge $\{u,v\}$. Players $P_{u}$ and $P_{v}$ check whether $u$ and $v$ were both matched, in which case no further action is required. Similarly, if neither of them was matched, players $P_{u}$ and $P_{v}$ simply add the edge $\{u,v\}$ to the matching. Note that this cannot cause any 3-augmenting paths, because this would require both $u$ and $v$ to have a free neighbor, contradicting the assumption that the matching prior to the update was maximal. The more challenging case is that only one of the two nodes, say $v$, was matched to some node $w$ whereas $u$ is free: Player $P_{v}$ instructs $P_{w}$ to check with all neighbors of $w$ whether $w$ has any free neighbor. This requires $P_{w}$ to send a single message to all players, and each player who has at least one free neighbor $x$ of $w$ (note that this is part of the players’ initial knowledge) responds affirmatively by sending $x$’s ID. If no such $x$ is found, we are again done, as it is straightforward to see that there cannot exist a 3-augmenting path. Otherwise, the four players $P_{u}$, $P_{v}$, $P_{w}$, and $P_{x}$ then coordinate to update the matching by replacing the edge $\{v,w\}$ with the two edges $\{u,v\}$ and $\{w,x\}$, which can be done in $O(1)$ additional rounds. Note that this results in all four involved vertices to become matched, and hence this cannot create any new 3-augmenting path in the resulting matching.
@@
      Suppose that $u$’s degree $d(u)$ is such that $d(u)\leqslant 2\sqrt{m}$. Then, $P_{u}$ uses Spreading (see Lem. 1) to broadcast the IDs of $u$’s neighbors, denoted by $N(u)$, to all players, which takes at most $O(\lceil\frac{\sqrt{m}}{k\beta}\rceil)$ rounds. As a result, every player $P_{w_{i}}$, who hosts a vertex $w_{i}$ that is matched to some $v_{i}\in N(u)$, knows that $v_{i}$ is matched to a neighbor of $u$. Next, $P_{w_{i}}$ checks with the other players whether $w_{i}$ has a free neighbor. We point out that $P_{w_{i}}$ may be hosting multiple vertices that are matched to (distinct) neighbors of $u$’s, and, in that case, $P_{w_{i}}$ performs this check via Spreading in parallel for all such vertices. By assumption, there are at most $2\sqrt{m}$ such matched vertices for which the players need to perform this check (in parallel), and hence Spreading takes $O(\lceil\frac{\sqrt{m}}{k\beta}\rceil)$ rounds to complete according to Lemma 1. Any player who hosts at least one free neighbor $x_{i}$ of $w_{i}$, picks one of them arbitrarily, and then broadcasts the ID of $x_{i}$; again, all of these (at most $2\sqrt{m}$) IDs are sent by the players via Spreading. Consequently, player $P_{u}$ learns at least one 3-augmenting path that $u$ is part of (if one exists). $P_{u}$ chooses one of these paths, say $(u,v_{i},w_{i},x_{i})$, and then directly coordinates with players $P_{v_{i}}$, $P_{w_{i}}$, and $P_{x_{i}}$ to replace the edge $\{v_{i},w_{i}\}$ with the edges $\{u,v_{i}\}$ and $\{w_{i},x_{i}\}$ in the matching. As this does not cause any vertices to become free that were matched before the edge deletion, this operation cannot create new 3-augmenting paths.
@@
      Figure 1. Case 2(b) in the proof of Theorem 1: The edge $\{u,v\}$ is deleted. Blue nodes and blue thick edges correspond to matched nodes and their matched edges. The grey-shaded and blue-shaded areas show the sets $Z$ and $Z^{\prime}$, respectively. Node $w^{\prime}$ is the low-degree neighbor found via sampling by $v^{\prime}$ and, in this example, removing the edge $\{w^{\prime},v\}$ from the matching will result in an additional 3-augmenting path that needs to be handled accordingly.
      Now, suppose that $d(u)>2\sqrt{m}$: Let $Z$ denote the set of nodes that are matched to $u$’s neighbors. By a simple counting argument, it follows that there is a subset $Z^{\prime}\subseteq Z$ of size at least $\frac{|Z|}{2}$, such that every $x\in Z^{\prime}$ has $d(x)\leqslant 2\sqrt{m}$. See Figure 1 for an example. We describe how $P_{u}$ can contact the players of $\Gamma=\Theta\left(\log n\right)$ randomly sampled vertices in $Z$: Note that $P_{u}$ may not know $Z$ (and hence cannot determine the players hosting the vertices in $Z$), but only knows the set of neighbors $N(u)$. To implement this step, $P_{u}$ randomly samples $\Gamma$ nodes in $N(u)$ and broadcasts the set of sampled node IDs, denoted by $S$, via Spreading to all players, which takes $O\left(\lceil\Gamma/\beta k\rceil\right)=O\left(\lceil\frac{\log n}{\beta k}\rceil\right)$ rounds. For each sampled $v\in S$ that is matched to some $w\in Z$, player $P_{v}$ asks $P_{w}$ to broadcast the ID of $w$ via Spreading. It follows that, with high probability, $P_{u}$ obtains the node ID of some low-degree vertex $w^{\prime}\in Z^{\prime}$ that is matched to some $v^{\prime}\in N(u)$. (To obtain a Las Vegas algorithm, we can simply repeat this step, if necessary.) Then, the players $P_{u}$, $P_{v^{\prime}}$, and $P_{w^{\prime}}$ add the edge $\{u,v^{\prime}\}$ to the matching and unmatch the edge $\{v^{\prime},w^{\prime}\}$ instead. As a result, there cannot exist any 3-augmenting path traversing $\{u,w\}$, as we assumed that $u$ does not have any free neighbors. However, this operation may have introduced 3-augmenting paths starting at the now-unmatched node $w^{\prime}$.
@@
Finally, we observe the players do not need to keep track of any state across updates apart from the matched edges, which ensures that the algorithm is also memoryless. ∎
@@ 4. Incremental $\frac{2}{3}$-approximate Maximum Matching under Batch-Dynamic Updates
In this section, we consider an adaptive adversary who can insert a batch of edges all at once.
From Lemma 1, we can see that, for any $\frac{2}{3}$-approximate maximum matching, there is no augmenting path with length $3$, which exists if and only if there is a matched edge whose both endpoints have free neighbors. Throughout this section, we restrict ourselves to incremental updates, i.e., batches of edge insertions. The following theorem states our main result for batch dynamic algorithms, which we prove in the remainder of this section.
@@ 4.1. Description and Analysis of the Algorithm
Our algorithm splits the batch of $\ell$ new edges into mini-batches of $\mathbf{b}=\lfloor\sqrt{k\beta}\rfloor$ edges each, and then it processes these mini-batches sequentially, such that the resulting matching after each mini-batch is maximal and does not contain 3-augmenting paths. Thus, we focus on describing the algorithm for just a single mini-batch $B$. Without loss of generality, we assume that no edges in $B$ connect two matched nodes; we show that the algorithm does not unmatch any currently matched vertices and hence we can simply ignore such edges if they do happen to exist.
We first introduce some notation. We use $M_{0}$ to denote the matching before the algorithm performs any updates for mini-batch $B$. We say that an edge is new if it is part of $B$, and old otherwise. Similarly, if an edge was inserted between $u$ and a free node $v$, we say that $v$ is a new free neighbor of $u$. Note that $v$ may be a new free neighbor for one node and simultaneously an old free neighbor for another.
The main algorithm consists of three conceptual phases:
@@
Initially, we want to ensure that all players know $B$, and thus we broadcast the entire set to all players using procedure Spreading (see Lem. 1). Throughout Phase 1, we focus on a subgraph $G_{1}$ that is small enough (i.e., consists of $O(\sqrt{k\beta})$ edges) to be collected and processed efficiently by just a single player. The edges forming $G_{1}$ consist of several subsets that we describe next:
@@
  • We will also process certain triangles in Phase 1 that may be part of 3-augmenting paths: To this end, consider any triangle $t=\{w,u,c\}$, for some vertices $u$, $w$, and $c$. We say that $t$ is a harmful triangle, if the edge $\{w,u\}$ is matched, $c$ is free, and at least one of $w$ or $u$ has an additional incident new edge to a free neighbor. Note that this includes the case that at least one of the other two unmatched edges of the triangle (i.e., $\{w,c\}$ or $\{u,c\}$) are new. Also note that there exists at most one triangle $\{c,w,u\}$ for a matched edge $\{w,u\}\in I_{1}$ before edge insertions (otherwise, there would have been a 3-augmenting path prior to processing $B$). The harmful triangles also contain such triangles. On the other hand, we say that $t$ is a harmless triangle if it is not harmful. Let $T$ be the set of edges forming harmful triangles, including all incident new free edges in $B$.
Figure 2. The graph $G_{1}$ used in Phase 1: Newly inserted edges are orange and matched edges are blue. A solid line indicates that the edge is in $G_{1}$, whereas a dashed line corresponds to a (possibly matched or new) edge that exists but is not included in $G_{1}$. In this example, $e_{1}\in F$, $e_{2}\in I_{1}$, $e_{3}\notin I_{1}$. Note that $t_{1}$ forms a harmful triangle and thus we include its edges and its incident new edges in $G_{1}$. On the other hand, $t_{2}$ is omitted, as it is a harmless triangle due to $e_{4}$ not being matched.
@@
Our analysis crucially relies on the invariant that there are no $3$-augmenting paths in subgraph $G_{1}$, which we guarantee by sending all of these edges to $P_{1}$, including the status of each edge (i.e., whether it is matched). Upon receiving $G_{1}$, player $P_{1}$ extends the current matching on $G_{1}$ to a (locally-computed) maximal matching $M^{\prime}$, fixes some ordering of the matched edges, and then iteratively removes any 3-augmenting path $p=(a,b,c,d)$, where $\{b,c\}$ is currently matched, by unmatching $\{b,c\}$ and instead matching $\{a,b\}$ and $\{c,d\}$. Subsequently, all other players update the state of their vertices and edges by learning the computed matching. Both of these communication steps are done by invoking Algorithm Spreading. We use $M_{1}$ to denote the updated matching on the whole graph.
@@
Assume towards a contradiction that there is an edge $e$ between two vertices in $V\cup X$. Since both endpoints of $e$ are free vertices at the start of Phase 2, and we do not unmatch any previously matched vertex in Phase 1 (see Lem. 3(a)), it follows that that the endpoints of $e$ were free also at the start of Phase 1. Thus, $e$ must be a new edge, as otherwise the matching that we started with was not maximal. Since $e\in F$, it must have been included in graph $G_{1}$, on which we have computed a maximal matching in Phase 1, which is a contradiction. ∎
@@
We first show the result for nodes in $W$. Assume towards a contradiction that $w_{i}\in W$ has a new free neighbor $v$. Recall that $w_{i}$ is matched to $u_{i}$ who also has a new free neighbor $v^{\prime}$. Moreover, by the definition of $I$, node $w_{i}$ also has an old free neighbor $x\in X$ and it holds that $x\neq v^{\prime}$. If $v=v^{\prime}$, then $(w_{i},u_{i},v^{\prime})$ forms a harmful triangle $t$, as there is a 3-augmenting path $p=(x,w_{i},u_{i},v^{\prime})$. By the definition of graph $G_{1}$ (see Lem. 2), we know that $p$ is part of $G_{1}$, contradicting Lemma 3(d). Now, suppose that $v\neq v^{\prime}$. It follows that there exists a 3-augmenting path $(v,w_{i},u_{i},v^{\prime})$ that is part of $\hat{I}_{1}\subseteq G_{1}$, because both $\{v,w_{i}\}$ as well as $\{u_{i},v\}$ are new free edges. This contradicts Lemma 3(c).
Now, assume towards a contradiction that some node $u_{i}\in U$ has a free old neighbor $x$. Let $x^{\prime}$ be the free old neighbor of $w_{i}$. In the case that $x\neq x^{\prime}$, there exists a 3-augmenting path $p=(x^{\prime},w_{i},u_{i},x)$. By Lemma 3(a) we did not unmatch any vertices in Phase 1, and thus it follows that $p$ also existed in the matching at the start of the update, which contradicts that the previous matching did not contain 3-augmenting paths. On the other hand, if $x=x^{\prime}$, we know that $u_{i}$ must also have a free new neighbor $v$. It follows that $(x,w_{i},u_{i})$ forms a harmful triangle, because there is a 3-augmenting path $(x,w_{i},u_{i},v)$, thus contradicting Lemma 3(d).
@@
Consider the matching $M_{1}$ resulting from Phase 1. Every 3-augmenting path in $G$ must traverse a matched edge in $I$.
@@
Suppose that there is a 3-augmenting path $(a,b,c,d)$, where the matched edge is $\{b,c\}\notin I$. It is not possible that both $a$ and $b$ have a new incident edge, because then $(a,b,c,d)$ would have been part of $\hat{I}_{1}\subseteq G_{1}$ and processed in Phase 1. Furthermore, it is also not possible that both $a$ and $b$ have an incident old free edge, as, in that case, the matching that we started with prior to the update had a 3-augmenting path, contradicting our assumption. Thus, it follows that either $\{a,b\}$ or $\{c,d\}$ must be a new edge, which allows us to conclude that $\{b,c\}\in I$. ∎
@@
We use the edge set $I$ (see Def. 4) to define a new bipartite graph $G_{X,W}$, on which we will operate in Phase 2: The vertices of $G_{X,W}$ consist of the sets $X$ and $W$. We include all edges from $G$ in $G_{X,W}$ that are of the form $\{x,w\}$, where $x\in X$ and $w\in W$, which ensures that $G_{X,W}$ is bipartite. In particular, we omit all edges between nodes in $W$ (recall that there cannot be any edges at this point between nodes in $X$ as all of them are free, see Lemma 5.) Figure 3(a) shows an example of $G_{X,W}$.
(a) Graph $G_{X,W}$
(b) Graph $G_{*}^{\prime}$
(c) Graph $G_{*}$
Figure 3. The subgraphs on which the algorithm operates in Phases 2 and 3. Thick blue edges are matched and orange edges are part of the batch of inserted edges. Free nodes are colored white.
Next, we make use of the result of (li2024dynamic), namely that there exists an algorithm that maintains a maximal matching on a graph with a small update time, when the adaptive adversary deletes a batch edges. A useful consequence of their result is that it yields a fast algorithm for computing a maximal matching on bipartite graphs, as we show next:
@@
Without loss of generality, assume that $|A|\leqslant|B|$. Since each player knows which of its vertices are in $A$ and which ones are in $B$, we can simply aggregate these counts at player $P_{1}$ and ensure that every player learns that $s=|A|$. Then, for each node $a\in A$, the player $P_{a}$ locally adds a neighbor $a^{\prime}$ to $H$ and pretends that the edge $\{a,a^{\prime}\}$ is matched, which ensures that there is a maximal matching on this extended graph. Note that the resulting graph partition is still balanced. Then, we remove all $\{a,a^{\prime}\}$ edges, to which the players respond by executing the update algorithm of (li2024dynamic). Since the number of deleted edges is $s$, Theorem 1.2 in (li2024dynamic) tells us that this algorithm completes in $O\left(\frac{s\log(n)}{\sqrt{\beta k}}\right)$ rounds, as required. ∎
@@
In the final phase, we operate on the subgraph $G_{*}^{\prime}=G[X^{\prime}\cup W\cup U\cup V]$. By definition, every edge in $I$ has at least one incident new edge (at the start of Phase 2), and the same is true for the vertices in $V$. This means that the number of vertices in $G_{*}^{\prime}$ is $O(\sqrt{k\beta})$, and thus the number of edges (new or old) in $G_{*}^{\prime}$ is $O({k\beta})$. We send all edges of $G_{*}^{\prime}$ to player $P_{1}$ in $O(1)$ rounds for local processing via Spreading (see Lem. 1). Starting from matching $M_{1}$, the local computation at $P_{1}$ produces the final matching $M$, by processing each $\{w_{i},u_{i}\}\in I$ in sequence, thereby iteratively augmenting the matching obtained in Phase 1. (Recall that, in Phase 2, we only computed a virtual matching without modifying the currently matched edges.) Note that we cannot simply instruct $P_{1}$ to locally compute a maximum matching on $G_{*}^{\prime}$, as we require the resulting matching to have certain properties, which we formally state in Lemma 10 below.
@@
For each edge $\{w_{i},u_{i}\}\in I$ ($i\geqslant 1$), player $P_{1}$ checks whether there is a 3-augmenting path $p=(x_{i},w_{i},u_{i},v)$ in $G_{*}^{\prime}$, for any free $v\in V$, where $x_{i}\in X^{\prime}$ denotes the vertex to which $w_{i}$ was matched in $M_{X,W}$, if it exists. If there is such a path, then $P_{1}$ updates $M$ by replacing $\{w_{i},u_{i}\}$ with the edges $\{x_{i},w_{i}\}$ and $\{u_{i},v\}$ in the matching, before moving on to edge $\{w_{i+1},u_{i+1}\}$.
@@
Next, player $P_{1}$ inspects every (currently) matched edge $\{w_{i},u_{i}\}\in I$ and checks whether there exists any 3-augmenting path for any free $x\in X^{\prime}$ and $v\in V$; note that it is not necessarily the case that $x=x_{i}$. If such a path exists, then we again resolve it by replacing $\{w_{i},u_{i}\}$ with the two other incident edges in the 3-augmenting path. After all edges in $I$ have been processed in the above manner, $P_{1}$ sends the final matching $M$ to every player via Spreading.
For our analysis, we define graph $G_{*}=G[X\cup W\cup U\cup V]$. Note that $G_{*}\supseteq G_{*}^{\prime}$, and $G_{*}$ may in fact contain up to $\Theta\left(n\right)$ nodes, which prohibits us from sending the entire graph to just a single player.